# Local Mechanisms of Compositional Generalization in Conditional Diffusion

### v2

@@ Appendix F Pixel- and Conditional-Locality Gradient Analyses
For Figure 7 we used the SDXL model (stabilityai/stable-diffusion-xl-base-1.0) as above. We used the following prompts for the feature-space analysis (listed with the shorthand used in the figure):
  • dog: “A dog, full body, highly detailed photograph.”
@@
  • horse: “A horse, full body, highly detailed photograph.”
  • cat: “A cat, full body, highly detailed photograph.”
  • vangogh: “An oil-painting in the style of Van Gogh.”
  • monet: “A watercolor-painting in the style of Monet.”
  • hokusai: “A woodblock print in the style of Hokusai.”
@@
  • sushi: “Eating sushi with chopsticks, highly detailed photograph.”
  • croissant: “Eating a croissant, highly detailed photograph.”
  • vangogh+cat+sushi: “A cat eating sushi with chopsticks, oil-painting in the style of Van Gogh.”
  • vangogh+cat: “A cat, full body, oil-painting in the style of Van Gogh.”
  • sushi+cat: “A cat eating sushi with chopsticks.”
  • unconditional: “”,
We used the attention_map_diffusers library wooyeolbaek (2025) to hook into SDXL’s cross-attention layers within the Down, Mid, and Up blocks. We ran with 15 inference steps total with the standard EulerDiscreteScheduler but extracted attention maps at intermediate times (step 1, 7, and 15 for low, mid, and high noise). To approximate the mean difference vectors for heuristic 3 we generated 10 samples per prompt and averaged the cross-attention maps, subtracting the unconditional average from each prompt average. Importantly, for each prompt, we actually selected and averaged only the specific cross-attention map corresponding to the single token most relevant to the concept (namely: dog, horse, cat, gogh, monet, sai, sushi, croissant) (some names get split into multiple tokens, so we choose the last one), and for the unconditional prompt we selected the end-of-sequence token (which we found received most of the attention). We then computed cosine similarities between the mean difference vectors for each pair of prompts. Figure 13 shows a larger version of the cosine similarity heuristic for all cross-attention layers shown in Figure 7, as well as the mid block activations at low and mid noise levels. We identify a possible structured feature-space within the first 5 cross-attention layers of the Mid block; note that this an empirical and subjective identification, but may help explain the model’s observed compositional success (see discussion in Section E.4).
For the Compositional attention analysis, we use the composite prompts listed above and compute the cosine similarity between the mean difference vector for the compositional prompt (e.g. vangogh+cat+sushi) evaluated at a specific token (e.g. cat) and the mean difference vector for the single “cat” prompt, evaluated at the cat token.

### v3

@@ G.2 Feature-space disentanglement
For Figure 5 we used the SDXL model (stabilityai/stable-diffusion-xl-base-1.0) as above. We used the following 24 prompts for the feature-space analysis, organized into 4 semantic categories (with shorthands used in figures):
@@
  • Animals (8): dog, cat, horse, rabbit, elephant, owl, fish, butterfly – e.g. “A dog, full body, highly detailed photograph.”
  • Art styles (7): monet, vangogh, hokusai, rembrandt, picasso, warhol, pixel_art – e.g. “An oil-painting in the style of Van Gogh.”
@@
  • Foods (6): sushi, croissant, coffee, pizza, ramen, cake – e.g. “Eating sushi with chopsticks, highly detailed photograph.”
  • Accessories (3): red_hat, sunglasses, scarf – e.g. “Wearing a red hat.”
  • unconditional: “” (empty string baseline).
To measure the F-LCS disentanglement heuristic (Lemma 2), we extract hidden-state activations (post-FFN outputs) from each BasicTransformerBlock in the SDXL U-net using PyTorch forward hooks. We generate 10 samples per prompt using 15 inference steps with the standard EulerDiscreteScheduler. For each prompt and seed, we record hidden states at each timestep and bin them into high ($t>666$), mid ($333<t\leq 666$), and low ($t\leq 333$) noise levels. We then average features across seeds, subtract the unconditional baseline, flatten each layer’s activations, and compute pairwise cosine similarities between all prompt pairs. We summarize disentanglement via within-category (mean cosine similarity among prompts in the same category) vs. between-category (mean cosine similarity among prompts in different categories) metrics, and their ratio. As baselines, we also compute the same metrics in the VAE latent space (the U-net output latents) and in pixel space (after VAE decoding), to verify that disentanglement is stronger in the network’s internal feature-space than in output spaces.
Figure 16: Additional detail for Figure 5. $24\times 24$ cosine similarity heatmaps (F-LCS heuristic, Lemma 2) for all 24 prompts across hidden-state activations in the down, mid, and up blocks (at high noise) and output spaces (latent/VAE, pixel). Prompts are grouped by category: animals (dog, cat, horse, rabbit, elephant, owl, fish, butterfly), art styles (monet, vangogh, hokusai, rembrandt, picasso, warhol, pixel_art), foods (sushi, croissant, coffee, pizza, ramen, cake), and accessories (red_hat, sunglasses, scarf). Block-diagonal category structure with crisper boundaries is visible in feature-space (especially down and mid blocks), while output spaces show weaker or absent structure. Within the art styles category, sub-cluster structure is also visible, with impressionist styles (monet, vangogh, hokusai, rembrandt) grouping separately from modern styles (picasso, warhol, pixel_art).
Appendix H Testing OOD prompts with a model trained on Flickr
Since the train sets of many large-scale text-to-image models like SDXL are not publicly known, the extent to which they are truly capable of OOD generalization is unclear. We therefore study a model for which the training set is known: Gu et al. (2023a)’s Matryoshka Diffusion Model (MDM) trained on subset of 50M Flickr images. In order to study OOD generalization, we designed candidate prompts and searched through the training set captions for keyword matches. We found some evidence of OOD composition on prompts with no conceptually-similar counterpart in the training set. Notably, we found that “a cat eating sushi” actually does appear in the train set; [^4] A train-set caption conceptually matching cat+sushi is: “A room with a wall painted with a mural of a cat eating sushi. The wall has a banner at the top with the words “DIADEMANG” written on it. The room has two low tables with cushions on the floor, and plates of sushi are placed on the tables. The lighting in the room is dim.” however, other similar prompts such as “a dog eating a croissant” or “a horse eating sushi” do not appear. One OOD example was “a cow jumping over a candlestick”: no conceptually similar prompts were found in the train set [^5] Two train-set captions matching the keywords “cow” and “candlestick” were found, for example: “A page from an old book with various crests, including one with scissors and a scissor-like symbol, one with a cow, one with a candlestick, one with a statue of a man holding a heart, and one with a shield with a cross and the words ‘La Comte Des Marechaux.’…” but neither conceptually represented a cow jumping over a candlestick., and yet the model is able to produce some plausible samples.
In Figure 17 we show several OOD example generations from MDM, as well as a feature-space analysis; we compare this to SDXL (although it is not known if the prompts are OOD for SDXL). We used the following prompts: “a cow jumping over a candlestick,” “a watercolor painting of a cow jumping over a candlestick,” “a dog eating a croissant,” “a watercolor painting of a dog eating a croissant,” “an oil-painting of an octopus flying through outer space,”, “an oil-painting of cat eating sushi with chopsticks.” (Only the last prompt is in-distribution for the Flickr training set; for all other compositional prompt a keyword search of the train set found no conceptual matches). MDM is evidently capable of some degree of compositional generalization, generating at least some plausible samples for the OOD prompts. The quality and prompt-fidelity of SDXL is higher, though we do not know whether the prompts are OOD for SDXL.
Figure 17 also shows feature-space cosine similarity experiments. We measure the Lemma 2 heuristic on hidden-state activations (post-FFN outputs of each SelfAttention module) in the innermost U-net of MDM, using 12 single-concept prompts spanning 6 semantic categories (4 animals, 2 art styles, 2 actions, 1 setting, 1 object, 2 foods), with 10 seeds per prompt and a single forward pass at high noise ($t=999$). We subtract the unconditional baseline before computing pairwise cosine similarity. The results show clear categorical disentanglement: within-category cosine similarity substantially exceeds between-category similarity across all blocks. For comparison, we run the same prompts through SDXL and find a qualitatively similar pattern at high noise, with even larger within/between gaps. These experiments offer preliminary – though far from conclusive – evidence that OOD compositional generalization is actually possible, and that compositional structure in feature-space may support it.
Figure 17: MDM generalization on OOD prompts. (Top) Sample generations from MDM (left) and SDXL (right) for compositional prompts. Rows are ordered from OOD (top) to in-distribution (bottom) for the MDM Flickr training set; OOD status for SDXL is unknown. (Middle right) Cosine similarity heatmaps of hidden-state activations (mid blocks, high noise) for SDXL and MDM, showing block-diagonal category structure (animals, styles, actions, foods). (Bottom) Within- vs. between-category cosine similarity (F-LCS heuristic, Lemma 2) across down, mid, and up blocks for both MDM ($t=999$) and SDXL (high noise), and a comparison of their disentanglement ratios. Both models show clear categorical disentanglement, with within-category similarity substantially exceeding between-category similarity across all blocks.